OpenStudy (anonymous):

Find the third degree polynomial with real coefficients such that 5 and are zeros and Answer X^3+ 5x^2 + 2x – 96 4x^3 - 20x^2– 4x – 80 x^3 – 5x^2+ 4x – 20 2x^3 – 10x^2 – 8x + 40

jimthompson5910 (jim_thompson5910):

"such that 5 and are zeros and"...that's just an odd phrase...

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

typo maybe?

OpenStudy (anonymous):

2i and f(2)=-96 sorry

jimthompson5910 (jim_thompson5910):

so it's Find the third degree polynomial with real coefficients such that 5 and 2i are zeros and f(2) = -96 ???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

If x = 2i is a zero, then x = -2i is also a zero since complex zeros come in conjugate pairs

jimthompson5910 (jim_thompson5910):

So we have 3 zeros: 5, 2i, -2i

jimthompson5910 (jim_thompson5910):

If x = 2i and x = -2i are zeros, then x = 2i or x = 2i x^2 = (2i)^2 or x^2 = (-2i)^2 x^2 = 4i^2 or x^2 = 4i^2 x^2 = 4i^2 x^2 = 4(-1) x^2 = -4 x^2 + 4 = 0 This means that x^2 + 4 is a factor

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next??

jimthompson5910 (jim_thompson5910):

Well if 5 is a zero, then x-5 is a factor

jimthompson5910 (jim_thompson5910):

So we have two factors x-5 and x^2+4

jimthompson5910 (jim_thompson5910):

So we then get k(x-5)(x^2+4)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next??

jimthompson5910 (jim_thompson5910):

you're told that f(2)=-96, so f(x) = k(x-5)(x^2+4) f(2) = k(2-5)((2)^2+4) -96 = k(2-5)((2)^2+4) solve for k

OpenStudy (anonymous):

\[ f(x)=b (x^2 + 4) (x - 5)\\ f[2]= -92\\ -24b=-96\\ b=4 \]

OpenStudy (anonymous):

wait so which equationis the answer??

jimthompson5910 (jim_thompson5910):

both are the same, we just used different variables up front

jimthompson5910 (jim_thompson5910):

and the order of factors is different (which doesn't matter)

OpenStudy (anonymous):

but my question is multiple choice the choices are above within the question

OpenStudy (anonymous):

It should be the second one if 4x is 16 x

OpenStudy (anonymous):

the 4x^3 one??

OpenStudy (anonymous):

\[ 4 (x-5) \left(x^2+4\right)=4 x^3-20 x^2+16 x-80 \]

OpenStudy (anonymous):

None of your choices are good.

OpenStudy (anonymous):

Are you sure you wrote them down right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The answer is none of the above

OpenStudy (anonymous):

ok thank you could you also help me with this other one find all the real zeros : f(x)=x^4-2x^3+10x^2-18x+9 f(x)=3x^4+27x^2+9